# Polarization of reflected X-ray emission from Sgr A molecular complex: multiple flares, multiple sources?

### v1

@@ 3 Data
4 X-ray images, light curves and spectra: epoch-to-epoch variations from Chandra
@@
The short flare reflection scenario is well established and tested on the data of extensive Chandra and XMM-Newton monitoring campaigns (e.g., Ponti et al., 2010; Clavel et al., 2013; Churazov et al., 2017c; Terrier et al., 2018; Ponti et al., 2013). Here, we are interested in the quantitative characterization of the reflection signal from several selected regions in order to use this as a basis for the spectropolarimetric analysis of the IXPE data.
@@ 4.1 Imaging
Due to the expected spectral shape of the reflected emission (Sunyaev & Churazov, 1996; Churazov et al., 2017a) and high foreground interstellar absorption towards the Galactic Center, the optimal band is commonly selected as 3–8 keV. This band contains both the reflection continuum and the most prominent fluorescent line of neutral (or only weakly ionized) iron at 6.4 keV. Also, it provides good leverage for spectral de-composition analysis needed to single out the ‘hot plasma’ ( keV) component (bulk of which is produced by unresolved population of accreting white dwarfs, so called Galactic ridge and bulge emission, Revnivtsev et al., 2009), as well an intermediate temperature component (2–3 keV), which is especially bright in the direction of the Sgr A complex. Finally, this band is almost insensitive to the dramatic decrease in Chandra’s response efficiency at low energy due to the build-up of the absorbing contaminant layer, allowing one a simple comparison between the epochs just at the band-integrated maps level (e.g., Churazov et al., 2017c).
@@ 4.2 Light curves
Epoch-to-epoch light curves of the emission in the 3–8 keV band for several regions of this field are shown in Fig. 3. One can see that the average surface brightness of the large circular region encompassing the Sgr A complex (black line) from 2000 to 2024 is relatively stable, even though the morphology of the emission within it was changing. Slight ( level) variations are present, some of which are caused by different positioning of this relatively large region in Chandra’s field of view, resulting in different weighting of various parts of it due to spatially-dependent vignetting and locations of the chip gaps, as well as the contribution of time-variable and transient point sources. This is illustrated by the presence of similar amplitude variation in a background region, which appears to contain a much smaller amount of reflected emission.
The other three regions are smaller in size and show substantial variations in surface brightness, with the brightest central part of the Sgr A complex demonstrating a rapid rise from 2013 to 2015 and then a decay after 2023. Decay of the diffuse emission surrounding the Arches cluster is well known and is also (partially) attributed to the decay of the reflected emission (e.g., Capelli et al., 2011; Krivonos et al., 2014; Clavel et al., 2014; Krivonos et al., 2017; Kuznetsova et al., 2019). More surprising is a decrease in surface brightness from the region dominated by the non-thermal emission from G0.13-0.11 nebula. However, spectral analysis of the variable part of the diffuse emission shows that it is also the decay of the reflection emission that drives this apparent trend (see Sect. 4.3).
@@ 4.3 Spectra
Spectroscopy of the time variable emission is a very powerful tool that allows us to get an idea about the spectral shape of the variable and non-variable components, which can then be used for single epoch spectral decomposition. In order to make this method most efficient, we select two epochs of observations that correspond to the minimal and maximal surface brightness in the regions of interest. We focus on three regions well within the Sgr A complex shown in Fig. 4 and the difference between observations taken in 2004 and 2022.
@@
We extract spectra in each of them and fit simultaneously by a model (described and exploited for Sgr A complex in Marin et al., 2023), allowing only the reflection component (derived from CREFL16 model555https://wwwmpa.mpa-garching.mpg.de/~churazov/crefl by Churazov et al., 2017a) to vary. The other components (‘the hot ( keV) plasma’ and ‘warm ( keV) plasma’) are essentially phenomenological descriptions of the temporally constant and spatially smooth emission of background and foreground sources (see the model definitions in Marin et al. 2023). Spectral analysis is performed using XSPEC666https://heasarc.gsfc.nasa.gov/xanadu/xspec/ software package (Arnaud, 1996, V12.14).
@@
Very good quality of the fits (=685 and 705 for 705 degrees of freedom for the ‘bg’ and ‘re’ reflection-dominated regions, and =740 with 705 degrees of freedom for the G0.13-0.11:‘nt’ region due to the presence of a non-thermal powerlaw component in this case) proves that the variable emission has the spectral shape that is indeed fully compatible with the reflection scenario (as shown in Fig. 5). In particular, the difference spectrum contains a hard X-ray continuum and a prominent fluorescent iron line that are treated separately during the fit to allow more flexibility in the spectral shape. The resulting non-reflection emission can be compared with the prediction of the spatial-decomposition model, which is capable of estimating the brightness of the bulge and ridge emission based on the projected stellar mass maps (Anastasopoulou et al., 2025). As a proxy for the spectral shape of this emission, we use spectra extracted from a deep observation of the bulge region (Revnivtsev et al., 2009) which is believed to contain negligible contribution of the diffuse reflected emission (even though featuring the 6.4 keV line due to the presence of such a line in spectra of certain types of accreting white dwarfs, see, e.g., Anastasopoulou et al. 2025). One can see that the emission predicted by this model does not exceed the derived non-variable part of the spectrum for these regions, ensuring that there is no overestimation of the purely reflection component.
@@
We can also see that even for the region dominated by non-thermal powerlaw emission of the G0.13-0.11 nebula, the spectrum of the emission responsible for the drop in brightness since 2004 is fully consistent with the reflection model. Interestingly, the relative strength, or equivalent width, of the fluorescent line differs slightly from region to region. Although this might be indication of a different reflection geometry, it is also within expected variations due to inhomogeneities in iron abundance or relative contribution of second scatterings (it is expected to be larger for fainter regions and feature a relatively stronger fluorescent line with a Compton shoulder profile)777Once the primary photons leave the cloud, it fades, but the doubly and multiply-scattered photons might still be detected with deep observations ( see, e.g. Sunyaev & Churazov, 1996; Churazov et al., 2002; Molaro et al., 2016; Kuznetsova et al., 2022; Khabibullin et al., 2022).. In the case of abundance variations, the predicted polarization signal is unchanged, while the other two cases predict correlated change in the polarization degree ( see, e.g., Churazov et al., 2002; Khabibullin et al., 2020a). In both of them, stronger line emission should correspond to a smaller polarization degree of the continuum, either as a result of contamination by less polarized second scattering emission or due to the phase function of the Compton scattering (for the spectral band of interest here, Thomson limit of the Compton scattering provides a fully adequate description) as a function of the cosine of the scattering angle. Another cause of spatially variable equivalent width of the fluorescent line might be intrinsic polarization of the primary emission, which in principle allows one to produce a high polarization degree in the continuum and a relatively strong emission line, but only at the cost of a relatively low continuum emission overall (Khabibullin et al., 2020a). In this case, the direction of the polarization vector might be very different from the unpolarized primary source prediction.
@@ 5.1 IXPE Stokes images and spectra
Figure 6 shows a comparison of Chandra and IXPE 3–8 keV images for the region of interest. The Chandra image corresponds to the middle 2022 epoch, while the IXPE image combines both 2022 and 2023 observation epochs. Only small morphological variations between 2022 and 2023 observations are expected from a comparison of the Chandra images taken in these epochs. The consistency of Chandra and IXPE images is very good (see Fig. 6), both in the large-scale diffuse emission and the three brightest individual regions – the Arches cluster, G0.13-0.11 PWN, and the brightest reflection spot in Sgr A complex. Some differences in morphology or relative brightness of certain structures are expected, given that the energy dependencies of the effective area of the two telescopes are different, resulting in slightly different weighting of various emission components in the combined 3–8 keV image. In what follows, we consider the combined IXPE dataset and demonstrate consistency of differences between the epochs with the expected statistical deviation level in Appendix A.
The consistency of the IXPE Stokes and Chandra data can be further validated by comparing spectra for several regions of interest. Figure 7 shows the IXPE Stokes spectra for two mutually exclusive regions: a large (r=4.5) circle region encompassing Sgr A (also excluding the region around the G0.13-0.11 nebula) and the brightest reflection spot in its center. Both spectra are perfectly fit with the models derived from the quasi-simultaneous Chandra data with only a slight normalization correction (factor 0.85) needed, which is well within the possible margin of the effect introduced by proper response function calculations for extended sources with complicated morphological structure. This factor is propagated further in a self-consistent manner, so the polarization degree calculations are not affected by this.
@@ 5.2 IXPE Stokes and images and spectra
Linearity of the Stokes parameters allows one to produce maps in (optimal) energy bands and smooth or bin them in order to suppress statistical noise. In Fig. 8, we show Stokes and images obtained from the full IXPE data set in the 3–6 keV band, where the signal-to-noise ratio for the reflected continuum is expected to be the highest. One can see dominance of the positive and for the G0.13-0.11 region, close to zero and negative for the large Sgr A region, and marginally positive and for the brightest reflection region. Also, the level of fluctuations due to background noise (and its increasing trend towards the boundaries) is visible. The latter effect is well known and described, arising from imperfections of the photoelectron tracks reconstruction near the boundaries of the Field-of-View (e.g., Di Marco et al., 2023). Our choice of the extraction regions ensures that the possible impact of this effect is minimal, given that all regions are a) always located close to the centers of the detectors; b) are large enough to fully encompass diffuse structures which are bright in reflected emission (even taking into account effective PSF of IXPE). We try to quantify the possible amplitude of other systematic effects in the Appendix A.
The described picture is consistent with the previous findings for the large Sgr A region and the non-thermal emission from G0.13-0.11, but the result obtained for the brightest reflection region is very surprising. Indeed, the Stokes and pattern for it appears more similar to the G0.13-0.11 region than to the Sgr A complex, meaning the direction of polarization is almost perpendicular to the expected one.
@@
Since we expect only the continuum part of the reflected emission component to be polarized, we build a model (Marin et al., 2023; Churazov et al., 2024) that allows one to simultaneously fit Stokes and data, varying polarization degree (PD) and polarization direction (in terms of the polarization angle ) for a fixed spectral shape and normalization of the reflected continuum for each region (found consistently in Chandra and IXPE Stokes spectral analysis described in Sect.s 4.3 and 5.1). The fitting is done via a standard steppar procedure in in XSPEC which involves calculating statistics on an extensive grid of PD and values ranging from 0 to 100% (defined with respect to the spectral model of the reflection continuum) and from to 90 degrees (in IAU convention), respectively. Namely, the model is composed of two additive components with tied spectral shape (given by the spectrum of the reflected continuum) and normalization (set by PD) multiplied by (for Stokes ) and (for Stokes ) user-defined functions. The results of this procedure are shown in Fig. 9. As already hinted by the maps of the Stokes parameters Q&U, the signal from the large Sgr A region is characterized by negative and close to zero values, while for the brightest reflection spot, both and are marginally positive with large statistical noise.
@@ 5.3 Inferred polarization parameters
The inferred confidence regions for the large circle (excluding the G0.13-0.11 and the brightest central region) and the brightest central region are shown in Fig. 10 along with the expectations for the reflection scenario with Sgr A* being the primary source (which sets PA, i.e. in this plot) and a range of the scattering angles (which sets the PD). The contours mark levels of 2.3, 4.61, 9.21 with respect to the best-fit parameters corresponding to 68%, 90%, and 99% confidence levels for two fitted parameters.
The best fit values for the large region are and deg (quoted uncertainties correspond to confidence interval), fully consistent with the first measurement reported in Marin et al. (2023), who found and deg. The detection of non-zero polarization is at the level of 3.9 significance based on the combined dataset. A slight increase in the PD is associated with the exclusion of the brightest reflection region, which appears to have almost orthogonal polarization direction (if any, see below), even though the polarization parameters inferred for the large region with and without exclusion of the brightest reflection spot are consistent within the uncertainties. Since this exclusion is made based on intensity only, without any prior knowledge of its polarization properties, no intentional bias is expected to be introduced by this procedure.
@@
As already hinted by the Stokes maps, the polarization signal extracted from the brightest reflection region is rather different: the best fit value for the PA is deg and polarization degree , i.e., significance of the detection of non-zero polarization falls just short of the two sigma level. On the other hand, the same polarization parameters as inferred for the large Sgr A region can be excluded at more than the 99% confidence level, given that the best fit polarization directions are almost perpendicular to each other. The 99% upper limit on the PD for the PA fixed at the direction towards Sgr A* is (1 upper limit is ), meaning that a situation with Sgr A* being the primary source and the cloud being significantly more distant (larger ) is still possible.
Given that the brightest reflection region appears to be encompassed by more extended diffuse emission with different polarization, one may assume that there is a contribution of the large-scale emission to the extracted signal from the brightest reflection region. In such a case, one would need to subtract (or model) this contribution when fitting the polarization spectra. Using the signal from the large region as an estimate for the background for the brightest reflection region, the fitting procedure results in (i.e., 2.4) and deg. Naturally, this increase in significance is driven partially by the fact that polarization in the background region is almost perpendicular to the marginally detected polarization in the brightest reflection region. The 99% upper limit on the PD for the PA fixed at the direction towards Sgr A* is (1 upper limit is ) in this case.
Finally, one might expect the impact of systematic effects for the brightest region in the center of the Field-of-View. Analysis of the individual datasets (two observational epochs and 3 DUs) does not reveal exceptional outliers, although a significant detection is achieved only in the combined data for one of the detector units (DU1). For the other two detectors, the hypothesis of polarization signal being identical to the one measured from the large region is excluded at the confidence level of at least 63% and 74% confidence levels (see Table 3).

### v2

@@ 3.2 Chandra data
4 X-ray images, light curves, and spectra: Epoch-to-epoch variations from Chandra
@@
The reflection scenario for the Sgr A complex is well established and tested on the data of extensive Chandra and XMM-Newton monitoring campaigns (Ponti et al., 2010, 2013; Clavel et al., 2013; Churazov et al., 2017c; Terrier et al., 2018; Alboslani et al., 2025). Here, we are interested in the quantitative characterization of the reflection signal from several selected regions in order to use this as a basis for the spectropolarimetric analysis of the IXPE data.
Due to the expected spectral shape of the reflected emission (Sunyaev and Churazov, 1996; Churazov et al., 2017a) and high foreground interstellar absorption toward the Galactic Center, the optimal band is commonly selected as 3–8 keV. This band contains both the reflection continuum and the most prominent fluorescent line of neutral (or only weakly ionized) iron at 6.4 keV. Also, it provides good leverage for spectral decomposition analysis needed to single out the ‘hot plasma’ ( keV) component (the bulk of which is produced by the unresolved population of accreting white dwarfs, so called Galactic ridge and bulge emission, Revnivtsev et al., 2009), as well an intermediate temperature component (2–3 keV), which is especially bright in the direction of the Sgr A complex. Finally, this band is almost insensitive to the dramatic decrease in Chandra’s response efficiency at low energies due to the build-up of the absorbing contaminant layer, allowing one a simple comparison between the epochs just at the band-integrated maps level (e.g., Churazov et al., 2017c).
@@ 4.2 Light curves
Epoch-to-epoch light curves of the emission in the 3–8 keV band for several regions of this field are shown in Fig. 3 (exact definitions of the regions are listed in Table 10). One can see that the average surface brightness of the large circular region encompassing the Sgr A complex (black line) from 2000 to 2024 is relatively stable, even though the morphology of the emission within it was changing. Slight ( level) variations are present, some of which are caused by different positioning of this relatively large region in Chandra’s FoV, resulting in different weighting of various parts of it due to spatially dependent vignetting and locations of the chip gaps, as well as the contribution of time-variable and transient point sources. This is illustrated by the presence of similar amplitude variation in a background region, which appears to contain a much smaller amount of reflected emission.
The other three regions are smaller in size and show substantial variations in surface brightness, with the brightest central part of the Sgr A complex demonstrating a rapid rise from 2013 to 2015 and then a decay after 2023. Decay of the diffuse emission surrounding the Arches cluster is well known and is also (partially) attributed to the decay of the reflected emission (e.g., Capelli et al., 2011; Krivonos et al., 2014; Clavel et al., 2014; Krivonos et al., 2017; Kuznetsova et al., 2019). More surprising is a decrease in surface brightness from the region dominated by the nonthermal emission from G0.13-0.11 nebula. However, spectral analysis of the variable part of the diffuse emission shows that it is also the decay of the reflection emission that drives this apparent trend (see Sect. 4.3).
@@ 4.3 Spectra
Spectroscopy of the time variable emission is a very powerful tool that allows us to get an idea about the spectral shape of the variable and non-variable components, which can then be used for single epoch spectral decomposition. In order to make this method most efficient, one can select two epochs of observations that correspond to the minimal and maximal 38 keV surface brightness in the regions of interest. We focus on three regions well within the Sgr A complex shown in Fig. 4 and consider the difference between observations taken in 2004 and 2022, optimizing for the brightest central region in Sgr A (see Fig. 3).
@@
We extract spectra in each of them and fit simultaneously by a model (described and exploited for Sgr A complex in Marin et al., 2023), allowing only the reflection component (derived from CREFL16 model666https://wwwmpa.mpa-garching.mpg.de/~churazov/crefl by Churazov et al., 2017a) to vary. The other components (‘the hot ( keV) plasma’ and ‘warm ( keV) plasma’) are essentially phenomenological descriptions of the temporally constant and spatially smooth emission of background and foreground sources (see the model definitions in Marin et al. 2023). The spectral shapes of all components were fixed (according to the parameters found by analysis of sensitive epoch-resolved XMM-Newton and Chandra data of Sgr A performed in Marin et al. 2023) and only their normalizations were fitted using standard procedures of XSPEC777https://heasarc.gsfc.nasa.gov/xanadu/xspec/ software package (Arnaud, 1996, V12.14). Although the spectral shape of both the reflection signal and the contaminating emission might vary from region to region, the phenomenological model constructed for the full Sgr A complex is flexible enough to describe these variations by changing the relative contributions of its components. In particular, in the case of the reflection signal, the CREFL16 model is split into continuum (CREFLcont) and fluorescent line (CREFLfluor) components, and freedom in their relative normalization is capable of absorbing possible variations in the spectral shape of the reflected emission given limited energy range of the available spectra. The non-reflection components of the model do not play any role in the polarization analysis, given that their spectral shapes are sufficiently ‘orthogonal’ to the spectral shape of the time-variable reflection components.
@@
Very good quality of the fits (=685 and 705 for 705 degrees of freedom for the ”bg” and ”re” reflection-dominated regions, and =740 with 705 degrees of freedom for the G0.13-0.11:”nt” region due to the presence of a nonthermal powerlaw component in this case) proves that the variable emission has the spectral shape that is indeed fully compatible with the reflection scenario (as shown in Fig. 5). In particular, the difference spectrum contains a hard X-ray continuum and a prominent fluorescent iron line that are treated separately during the fit to allow more flexibility in the spectral shape. The resulting non-reflection emission can be compared with the prediction of the spatial-decomposition model, which is capable of estimating the brightness of the bulge and ridge emission based on the projected stellar mass maps (Anastasopoulou et al., 2025). As a proxy for the spectral shape of this emission, we use spectra extracted from a deep observation of the bulge region (Revnivtsev et al., 2009) which is believed to contain negligible contribution of the diffuse reflected emission (even though featuring the 6.4 keV line due to the presence of such a line in spectra of certain types of accreting white dwarfs; see, e.g., Anastasopoulou et al. 2025). One can see that the emission predicted by this model does not exceed the derived non-variable part of the spectrum for these regions, which we consider is a useful sanity check for the time variability-informed spectral decomposition.
@@
We also observed that even for the region dominated by nonthermal power-law emission of the G0.13-0.11 nebula, the spectrum of the emission responsible for the drop in brightness since 2004 is fully consistent with the reflection model. Interestingly, the relative strength, or equivalent width, of the fluorescent line differs slightly from region to region. Although this might be an indication of a different reflection geometry, it is also within expected variations due to inhomogeneities in iron abundance or the relative contribution of second scatterings (which is expected to be larger for fainter regions and feature a stronger fluorescent line with a Compton shoulder profile).888Once the primary photons leave the cloud, the cloud fades, but the doubly and multiply scattered photons might still be detected with deep observations ( see, e.g., Sunyaev and Churazov, 1996; Churazov et al., 2002; Molaro et al., 2016; Kuznetsova et al., 2022; Khabibullin et al., 2022). In the case of abundance variations, the predicted polarization signal is unchanged, while the other two cases predict correlated change in the polarization degree ( see, e.g., Churazov et al., 2002; Khabibullin et al., 2020b). In these two cases, stronger line emission should correspond to a smaller polarization degree of the continuum, either as a result of contamination by less polarized second scattering emission or due to the phase function of the Compton scattering (for the spectral band of interest here, Thomson limit of the Compton scattering provides a fully adequate description) as a function of the cosine of the scattering angle. Another cause of spatially variable equivalent width of the fluorescent line might be intrinsic polarization of the primary emission, which in principle allows one to produce a high polarization degree in the continuum and a relatively strong emission line, but only at the cost of a relatively low continuum emission overall (Khabibullin et al., 2020b). In this case, the direction of the polarization vector might be very different from the unpolarized primary source prediction.
@@ 5.1 IXPE Stokes images and spectra
Figure 6 shows a comparison of Chandra and IXPE 3–8 keV images for the region of interest. The Chandra image corresponds to the middle 2022 epoch, while the IXPE image combines both 2022 and 2023 observation epochs. Only small morphological variations between 2022 and 2023 observations are expected from a comparison of the Chandra images taken in these epochs. The consistency of Chandra and IXPE images is very good (see Fig. 6), both in the large-scale diffuse emission and for the three brightest individual regions – the Arches cluster, the G0.13-0.11 PWN, and the brightest reflection spot in the Sgr A complex. Some differences in morphology or relative brightness of certain structures are expected, given that the energy dependencies of the effective area of the two telescopes are different, resulting in slightly different weighting of various emission components in the combined 3–8 keV image. In what follows, we consider the combined IXPE dataset and demonstrate consistency of differences between the epochs with the expected statistical deviation level in Appendix B.
The consistency of the IXPE Stokes and Chandra data can be further validated by comparing solid angle-normalized spectra for several regions of interest. Figure 7 shows the IXPE Stokes spectra (corresponding to the surface brightness per arcmin2) for two mutually exclusive regions: a large (r=4.5) circle region encompassing Sgr A (also excluding the region around the G0.13-0.11 nebula) and the brightest reflection spot in its center. Both area-normalized spectra are perfectly fit with the models derived from the quasi-simultaneous Chandra data with only a slight normalization correction (factor 0.85) needed, which is well within the possible margin of the effect introduced by proper response function calculations for extended sources with complicated morphological structure. This factor is propagated further in a self-consistent manner, so the polarization degree calculations are not affected by this.
@@ 5.2 IXPE Stokes and images and spectra
Linearity of the Stokes parameters allows one to produce maps in (optimal) energy bands and smooth or bin them in order to suppress statistical noise. In Fig. 8, we show Stokes and images obtained from the full IXPE dataset in the 3–6 keV band, where the signal-to-noise ratio for the reflected continuum is expected to be the highest. One can see dominance of the positive and for the G0.13-0.11 region, close to zero and negative for the large Sgr A region, and marginally positive and for the brightest reflection region. Also, the level of fluctuations due to background noise (and its increasing trend toward the boundaries) is visible. The latter effect is well known and described, arising from imperfections of the photoelectron tracks reconstruction near the boundaries of the FoV (e.g., Di Marco et al., 2023). Our choice of the extraction regions ensures that the possible impact of this effect is minimal, given that all regions are a) always located close to the centers of the detectors; b) are large enough to fully encompass diffuse structures which are bright in reflected emission (even taking into account the effective PSF of IXPE). We try to quantify the possible amplitude of other systematic effects in Appendix B.
The described picture is consistent with the previous findings for the large Sgr A region and the nonthermal emission from G0.13-0.11, but the result obtained for the brightest reflection region is very surprising. Indeed, the Stokes and pattern for it appears more similar to the G0.13-0.11 region than to the Sgr A complex, meaning the direction of polarization is almost perpendicular to the expected one.
@@
Since we expect only the continuum part of the reflected emission component to be polarized, we build a model (Marin et al., 2023; Churazov et al., 2024) that allows one to simultaneously fit Stokes and data, varying polarization degree (PD) and polarization direction (in terms of the PA ) for a fixed spectral shape and normalization of the reflected continuum for each region (found consistently in Chandra and IXPE Stokes spectral analysis described in Sect. 4.3 and 5.1). The fitting is done via a standard steppar procedure in in XSPEC which involves calculating statistics on an extensive grid of PD and values ranging from 0 to 100% (defined with respect to the spectral model of the reflection continuum) and from to 90 degrees (in IAU convention), respectively. Namely, the model is composed of two additive components with tied spectral shape (given by the spectrum of the reflected continuum) and normalization (set by PD) multiplied by (for Stokes ) and (for Stokes ) user-defined functions. The results of this procedure are shown in Fig. 9. As already hinted by the maps of the Stokes parameters Q&U, the signal from the large Sgr A region is characterized by negative and close to zero values, while for the brightest reflection spot, both and are marginally positive with large statistical noise.
@@ 5.3 Inferred polarization parameters
The inferred confidence regions for the large circle (excluding the G0.13-0.11 and the brightest central region) and the brightest central region are shown in Fig. 10 along with the expectations for the reflection scenario with Sgr A* being the primary source (which sets PA, i.e., in this plot) and a range of the scattering angles (which sets the PD). The contours mark levels of 2.3, 4.61, 9.21 with respect to the best-fit parameters corresponding to 68%, 90%, and 99% confidence levels for two fitted parameters.
The best fit values for the large region are and deg (quoted uncertainties correspond to confidence interval), fully consistent with the first measurement reported in Marin et al. (2023), who found and deg. The detection of nonzero polarization is at the level of 3.9 significance based on the combined dataset. A slight increase in the PD is associated with the exclusion of the brightest reflection region, which appears to have almost orthogonal polarization direction (if any; see below), even though the polarization parameters inferred for the large region with and without exclusion of the brightest reflection spot are consistent within the uncertainties we list polarization parameters measured for both cases in two different epochs in Appendix B for direct comparison. Since this exclusion is made based on intensity only, without any prior knowledge of its polarization properties, no intentional bias is expected to be introduced by this procedure.
@@
As already hinted by the Stokes maps, the polarization signal extracted from the brightest reflection region is rather different. The best fit value for the PA is deg and polarization degree . That is, the significance of the detection of nonzero polarization falls just short of the two sigma level. On the other hand, the same polarization parameters as inferred for the large Sgr A region can be excluded with a confidence level of more than 99%, given that the best-fit polarization directions are almost perpendicular to each other. The 99% upper limit on the PD for the PA fixed at the direction toward Sgr A* is (1 upper limit is ), meaning that a situation with Sgr A* as the primary source and the cloud as significantly more distant (larger ) is still possible.
Given that the brightest reflection region appears to be encompassed by more extended diffuse emission with different polarization, one may assume that there is a contribution of the large-scale emission to the extracted signal from the brightest reflection region. In such a case, one would need to subtract (or model) this contribution when fitting the polarization spectra. Using the signal from the large region as an estimate for the background for the brightest reflection region, the fitting procedure results in (i.e., 2.4) and deg. Naturally, this increase in significance is driven partially by the fact that polarization in the background region is almost perpendicular to the marginally detected polarization in the brightest reflection region. The 99% upper limit on the PD for the PA fixed at the direction toward Sgr A* is (1 upper limit is ) in this case.
Finally, one might expect the impact of systematic effects for the brightest region in the center of the FoV. Analysis of the individual datasets (two observational epochs and 3 DUs) does not reveal exceptional outliers, although a significant detection is achieved only in the combined data for one of the detector units (DU1). For the other two detectors, the hypothesis of polarization signal being identical to the one measured from the large region is excluded at the confidence level of at least 63% and 74% confidence levels (see Table 4).